OpenStudy (anonymous):

Sarah wants to hang a mirror in her room but the mirror and frame must not have an area larger than 15 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.)

OpenStudy (anonymous):

a. 3x2 + 14x − 2 = 0 b. 4x2 + 14x − 3 = 0 c. x2 + 12x − 14 = 0 d. 2x2 + 7x − 14 = 0

OpenStudy (anonymous):

For a,b,c,d are you suppose to factor

OpenStudy (anonymous):

@iamlegend I'm not sure, if it's easier, why not

OpenStudy (anonymous):

if it doesn't say i'm going to assume solving by factoring, quadratic formula, or solutions

OpenStudy (anonymous):

@iamlegend okay

OpenStudy (anonymous):

one sec, be right back

OpenStudy (anonymous):

ok, srry about that

OpenStudy (anonymous):

so for (a). you can use quadratic formula Step 1:- A = 1, B = 5 and C = -6 Step 2:- D = b2 -4ac, Step 3:- D= 5x2 – 4 * 6, Step 4:- D= 25 – 24, Step 5:- D = 1, since D>0, So x1 = (-2), x2 = (-3), The factors are (x + 2), (x + 3). Answer = (x + 2), (x + 3). understand?

OpenStudy (anonymous):

Use factoring or quadratic formula for b,c,d as well, and let me know if you need help.

OpenStudy (anonymous):

I think so, would it be C?

OpenStudy (anonymous):

would what be C? that was for problem (a)

OpenStudy (anonymous):

the answer would be c, right? c is x2 + 12x − 14 = 0

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

really?! thank you!

OpenStudy (anonymous):

your welcome